# Learning Contrastive Feature Representations for Facial Action Unit Detection

### v7

@@ 4.2.1 Evaluation on BP4D
Methods in the second group mainly utilize the semantic prior knowledge, including MMA-Net [32], SEV-Net [33] and ETDF [34]. Our approach demonstrates an average improvement of 2.7%, 2.2% and 1.0%, respectively. This phenomenon can be attributed to the inherent limitations of these methods in capturing the fine-grained and localized cues associated with individual AUs, as simple semantic descriptions often fail to reflect the underlying visual complexity. Additionally, their dependence on predefined semantic priors may introduce inductive biases that hinder the model’s ability to flexibly adapt to the nuanced and heterogeneous nature of facial behaviors across individuals and contexts.
Methods in the third group leverage attention mechanisms to selectively focus on regional facial areas surrounding AUs, including ARL [35], AC2[3] and ESCDA [5]. Our approach achieves a notable improvement, with an increase in F1-score ranging from 0.6% to 5.0% over these methods. This advancement is primarily due to the feature extraction methods of these frameworks neglecting critical correlation information among AUs and failing to specifically address the class imbalance issue for each AU. Furthermore, their reliance on attention mechanisms may lead to suboptimal encoding when attention is misdirected or insufficiently localized, limiting the precision of AU-specific feature representations.
Methods in the fourth group focus on capturing correlational information among AUs, including AAR [28], SACL [36] and MEGraph [22]. Our approach exhibits a performance advantage of approximately 0.5%-2.3% in F1-score. This underscores the efficacy of our supervised discriminative contrastive learning paradigm in effectively capturing differential information between AUs from each sample pair instead of pixel-level information of the entire face associated with AU. It is noteworthy that the encoder used in our method is only the same as the network MEGraph used for the first stage of training, with reduced parameters by nearly 7M. This highlights learning only the difference information among features is more efficient than learning the whole pixel-level information.
Methods in the last group are based on contrastive learning, including SIMCLR [11], MoCo [12], EmoCo [19], CLP [9], KSRL [13] and CLEF [14]. When compared to these methods, our approach demonstrates a performance advantage of approximately 0.2% to 16.3% in F1-score. This improvement is attributed to the efficacy of our proposed negative sample re-weighting strategy and positive sample sampling strategy, specifically tailored for the AU detection task using static images. In addition, compared to contrastive learning methods based on weighted cross-entropy loss, our approach significantly reduces the extraction of pixel-level redundant information.
@@ 4.2.2 Evaluation on DISFA
Comparable results are observed on the DISFA dataset (Table 3), where AUNCE achieves the highest overall AU detection performance, though the improvement margin is smaller than on BP4D (Table 2). This discrepancy can be attributed to the more severe class imbalance, annotation noise, and uneven AU distribution in DISFA, which collectively hinder model generalization. Nevertheless, unlike methods such as MEGraph [22], CLP [9], and SEV-Net [33] that perform well on BP4D but degrade noticeably on DISFA, AUNCE maintains stable and superior performance across both datasets, benefiting from the proposed negative sample re-weighting strategy that mitigates the effects of noisy and imbalanced samples and enhances robustness across diverse AUs.
@@ 4.2.6 Cross dataset validation of AUNCE
To assess the generalization capability and cross-dataset robustness of our framework, we compare AUNCE against three baselines: a model using the same encoder with weighted cross-entropy loss (WCE), AC2[3], and D2CA [20], the latter of which is specifically designed for cross-domain AU detection. The comparative results are summarized in Table 7. Given that 10 AUs in the GFT dataset overlap with 12 AUs in BP4D, we conduct cross-dataset evaluations between BP4D and GFT.
@@
MEGraph (Onestage) [22] 2022IJCAI WCE 63.6 60.8
@@ 4.3.1 Compared with weighted cross-entropy loss and weighted focal loss
The primary innovation of this paper is the introduction of AUNCE, a novel contrastive loss designed to be compatible with various AU detection encoders. To evaluate its effectiveness, we compare the proposed AUNCE loss against the traditional weighted cross-entropy loss (WCE) and the recently introduced weighted focal loss (WFL) [10] during the pretraining stage, using multiple AU detection encoders. While numerous loss functions have been proposed to address sample imbalance, WFL has emerged as a widely adopted and empirically validated approach for mitigating positive–negative imbalance in AU detection. The results, presented in Table 8, demonstrate that AUNCE consistently outperforms the other losses in AU detection. Notably, the MEGraph encoder achieves the highest performance, with scores of 66.1% and 66.4% on BP4D and DISFA datasets, respectively, when trained using AUNCE. This highlights the advantage of emphasizing the capture of distinct features from individual samples, rather than solely relying on pixel-level information of the entire face in relation to AU detection.
Table 9: Comparison of the computational complexity and efficiency of the weighted cross-entropy loss (WCE), the weighted focal loss (WFL), and the proposed AUNCE loss on BP4D and DISFA dataset, where encoders of these three losses are both MEGraph with the first-stage training network. Labels denote ”N = samples” and ”C = classes”.
@@ 4.3.2 Evaluating other components
We further assess the effectiveness of all proposed components. Quantitative results of ablation study are summarized in Table 10. The baseline (BL) is based on SIMCLR [11], with positive sample pairs replaced by images most similar to the original ones. Other components include the weight coefficient ($w_{i}$) for both WCE and AUNCE, the positive sample sampling strategy (PS), and the negative sample re-weighting strategy (NS).
@@
Figure 5: T-SNE visualization of feature representations on BP4D and DISFA datasets. Colors indicates whether AU1 exists. Green (#5BAB54) points denote negative samples (Label = 0), and purple (#885F9C) points denote positive samples (Label = 1). Top row: (a) Representations optimized by WCE on BP4D dataset, (b)-(f) Visualizations of ablation study on BP4D dataset, respectively represent model A-E. Bottom row: (g) Representations optimized by AUNCE on DISFA dataset, (h)-(l) Visualizations of ablation study on DISFA dataset, respectively represent model A-E.
@@ 4.4.1 $p_{1},p_{2},p_{3},p_{4}$, the probabilities of various positive sample pairs
The experimental results are presented in Table 11. When $p_{1}$=0.15, $p_{2}$=0.15, $p_{3}$=0.7, and $p_{4}$=0, the experimental result is optimal. The InfoMin [40] principle provides a theoretical explanation for the above results. The core of the InfoMin principle is to provide a criterion for judging the quality of a positive sample: a high-quality positive sample should retain the essential information required for the downstream task (AU detection) while minimizing its mutual information with the anchor as much as possible. Here, the ”anchor” refers to the original image or the core features of the target AU, and the ”positive examples” correspond to the samples represented by $p_{1}$, $p_{2}$, and $p_{3}$. For $p_{4}$-related samples, their mutual information with the anchor is extremely low due to their intrinsic noise nature, thus explaining why $p_{4}=0$ in the optimal configuration. The $p_{3}$-related samples (mixed positive samples) have moderate mutual information with the anchor, which balances the retention of positive class common features and the avoidance of information redundancy. As for $p_{1}$- and $p_{2}$-related samples, although they can filter noise, their mutual information with the anchor usually exceeds the ”sweet spot”. An increase in their proportion will lead to information redundancy and subsequent performance degradation, which justifies their optimal value of 0.15.
@@ 5 Conclusion
In this paper, we propose a discriminative contrastive learning framework tailored for AU detection, introducing a novel contrastive loss named AUNCE. Building upon the InfoNCE loss, AUNCE refines the task by shifting focus from learning pixel-level information of the entire face to capturing subtle feature variations during AU activation, thus improving detection efficiency. We also introduce a positive sample sampling strategy that constructs three types of positive sample pairs, selectively augmenting positive samples, especially in the presence of noisy and false AU labels. This approach ensures the model is trained on more reliable examples, improving upon conventional methods that lack adaptive sampling. To address class imbalance in AU detection, we implement a negative sample re-weighting strategy that prioritizes minority AUs, helping the model focus on hard-to-detect samples. This strategy promotes hard example mining by giving more weight to underrepresented classes, enhancing model performance and generalization. Experimental results demonstrate that AUNCE consistently outperforms existing methods, particularly those utilizing weighted cross-entropy loss. While current validation is limited to static-image datasets, it is worth noting that these datasets are derived from dynamic video sources. Hence, the proposed framework can be naturally extended to dynamic AU sequences.
Despite the proven effectiveness of our approach, hyperparameter selection emerges as a notable limitation. Model performance is sensitive to specific parameters, with optimal configurations varying across datasets—particularly the mixing ratio of positive samples. Its optimal value fluctuates significantly with changes in data quality and noise rates in real-world scenarios, yet the current determination of this ratio still relies on empirical adjustments. Such variability poses a constraint on reproducibility, given that the current parameter tuning process is largely empirical and lacks generalizability across diverse scenarios. Future work will first focus on developing adaptive optimization strategies—specifically designing dynamic adjustment mechanisms for the positive sample mixing ratio based on real-time data quality and noise levels—to enhance the model’s robustness, generalization, and reproducibility across tasks.
@@ Acknowledgement
This research was supported by the National Natural Science Foundation of China (Nos. 62176221, 61572407), and Sichuan Province Science and Technology Support Program (Nos. 2024NSFTD0036, 2024ZHCG0166).

### v9

@@ 4.2.1 Evaluation on BP4D
Methods in the second group mainly utilize the semantic prior knowledge, including MMA-Net [shang2023mma], SEV-Net [yang2021exploiting] and ETDF [chang2025facial]. Our approach demonstrates an average improvement of 2.7%, 2.2% and 1.0%, respectively. This phenomenon can be attributed to the inherent limitations of these methods in capturing the fine-grained and localized cues associated with individual AUs, as simple semantic descriptions often fail to reflect the underlying visual complexity. Additionally, their dependence on predefined semantic priors may introduce inductive biases that hinder the model’s ability to flexibly adapt to the nuanced and heterogeneous nature of facial behaviors across individuals and contexts.
Methods in the third group leverage attention mechanisms to selectively focus on regional facial areas surrounding AUs, including ARL [shao2022facial], AC2[shao2025facial] and ESCDA [shao2026constrained]. Our approach achieves a notable improvement, with an increase in F1-score ranging from 0.6% to 5.0% over these methods. This advancement is primarily due to the feature extraction methods of these frameworks neglecting critical correlation information among AUs and failing to specifically address the class imbalance issue for each AU. Furthermore, their reliance on attention mechanisms may lead to suboptimal encoding when attention is misdirected or insufficiently localized, limiting the precision of AU-specific feature representations.
Methods in the fourth group focus on capturing correlational information among AUs, including AAR [shao2023facial], SACL [liu2025multi] and MEGraph [luo2022learning]. Our approach exhibits a performance advantage of approximately 0.5%-2.3% in F1-score. This underscores the efficacy of our supervised discriminative contrastive learning paradigm in effectively capturing differential information between AUs from each sample pair instead of pixel-level information of the entire face associated with AU. It is noteworthy that the encoder used in our method is only the same as the network MEGraph used for the first stage of training, with reduced parameters by nearly 7M. This highlights learning only the difference information among features is more efficient than learning the whole pixel-level information.
Methods in the last group are based on contrastive learning, including SIMCLR [chen2020simple], MoCo [he2020momentum], EmoCo [sun2021emotion], CLP [li2024contrastive], KSRL [chang2022knowledge] and CLEF [zhang2023weakly]. When compared to these methods, our approach demonstrates a performance advantage of approximately 0.2% to 16.3% in F1-score. This improvement is attributed to the efficacy of our proposed negative sample re-weighting strategy and positive sample sampling strategy, specifically tailored for the AU detection task using static images. In addition, compared to contrastive learning methods based on weighted cross-entropy loss, our approach significantly reduces the extraction of pixel-level redundant information.
@@ 4.2.2 Evaluation on DISFA
Comparable outcomes are evident in the DISFA dataset, as outlined in Table 3, where AUNCE attains the highest overall performance in AU detection. Furthermore, a comparison between Table 2 and Table 3 reveals that recent methods such as EmoCo, CLP, and MEGraph demonstrate strong performance on the BP4D dataset but achieve only mediocre results on the DISFA dataset. This performance gap is likely attributed to the more pronounced class imbalance issue in DISFA, as illustrated in Table 1. In contrast, AUNCE consistently delivers robust performance on both BP4D and DISFA datasets and achieves more stable results across various AUs due to our proposed negative sample re-weighting strategy.
@@ 4.2.6 Cross dataset validation of AUNCE
To assess the generalization capability and cross-dataset robustness of our framework, we compare AUNCE against three baselines: a model using the same encoder with weighted cross-entropy loss (WCE), AC2[shao2025facial], and D2CA [li2025decoupled], the latter of which is specifically designed for cross-domain AU detection. The comparative results are summarized in Table 7. Given that 10 AUs in the GFT dataset overlap with 12 AUs in BP4D, we conduct cross-dataset evaluations between BP4D and GFT.
@@
MEGraph (Onestage) [luo2022learning] 2022IJCAI WCE 63.6 60.8
@@ 4.3.1 Compared with weighted cross-entropy loss and weighted focal loss
The primary innovation of this paper is the introduction of AUNCE, a novel contrastive loss designed to be compatible with various AU detection encoders. To evaluate its effectiveness, we compare the proposed AUNCE loss against the traditional weighted cross-entropy loss (WCE) and the recently introduced weighted focal loss (WFL) [shang2024facial] during the pretraining stage, using multiple AU detection encoders. While numerous loss functions have been proposed to address sample imbalance, WFL has emerged as a widely adopted and empirically validated approach for mitigating positive–negative imbalance in AU detection. The results, presented in Table 8, demonstrate that AUNCE consistently outperforms the other losses in AU detection. Notably, the MEGraph encoder achieves the highest performance, with scores of 66.1% and 66.4% on BP4D and DISFA datasets, respectively, when trained using AUNCE. This highlights the advantage of emphasizing the capture of distinct features from individual samples, rather than solely relying on pixel-level information of the entire face in relation to AU detection.
Table 9: Comparison of the computational complexity and efficiency of the weighted cross-entropy loss (WCE), the weighted focal loss (WFL), and the proposed AUNCE loss on BP4D and DISFA dataset, where encoders of these three losses are both MEGraph with the first-stage training network.
@@ 4.3.2 Evaluating other components
We further assess the effectiveness of all proposed components. Quantitative results of ablation study are summarized in Table 10. The baseline (BL) is based on SIMCLR [chen2020simple], with positive sample pairs replaced by images most similar to the original ones. Other components include the weight coefficient ($w_{i}$) for both WCE and AUNCE, the positive sample sampling strategy (PS), and the negative sample re-weighting strategy (NS).
@@
Figure 5: T-SNE visualization of feature representations on BP4D and DISFA datasets. Colors indicates whether AU1 exists. Top row: (a) Representations optimized by WCE on BP4D dataset, (b)-(f) Visualizations of ablation study on BP4D dataset, respectively represent model A-E. Bottom row: (g) Representations optimized by AUNCE on DISFA dataset, (h)-(l) Visualizations of ablation study on DISFA dataset, respectively represent model A-E.
@@ 4.3.3 $p_{1},p_{2},p_{3},p_{4}$, the probabilities of various positive sample pairs
The experimental results are presented in Table 11. When $p_{1}$=0.15, $p_{2}$=0.15, $p_{3}$=0.7, and $p_{4}$=0, the experimental result is optimal. It is observed that images with the lowest similarity to original images contribute minimally to the training process, as these images essentially represent noise and errors. The mixture of all positive samples plays a pivotal role, as it encourages the feature representation of each instance to be closer to the class centroid, facilitating the capture of essential characteristics of the positive class. The indispensability of the remaining two positive samples lies in their ability to mitigate the impact of noisy and false samples, thereby reinforcing the robustness of the training process.
@@ 5 Conclusion
In this paper, we propose a discriminative contrastive learning framework tailored for AU detection, introducing a novel contrastive loss named AUNCE. Building upon the InfoNCE loss, AUNCE refines the task by shifting focus from learning pixel-level information of the entire face to capturing subtle feature variations during AU activation, thus improving detection efficiency. We also introduce a positive sample sampling strategy that constructs three types of positive sample pairs, selectively augmenting positive samples, especially in the presence of noisy and false AU labels. This approach ensures the model is trained on more reliable examples, improving upon conventional methods that lack adaptive sampling. To address class imbalance in AU detection, we implement a negative sample re-weighting strategy that prioritizes minority AUs, helping the model focus on hard-to-detect samples. This strategy promotes hard example mining by giving more weight to underrepresented classes, enhancing model performance and generalization. Experimental results demonstrate that AUNCE consistently outperforms existing methods, particularly those utilizing weighted cross-entropy loss.
Despite the overall effectiveness of our approach, hyperparameter selection remains a limitation. Model performance is sensitive to certain parameters, with optimal configurations varying across datasets. The current tuning process is largely empirical and lacks generalizability. Future work will focus on developing adaptive hyperparameter optimization strategies to enhance robustness and generalization across diverse tasks. In addition, we plan to incorporate recently proposed loss functions designed to address class imbalance into the AU detection task. These include techniques based on class reweighting, margin adjustment, and adaptive sampling, which have shown promise in improving performance on underrepresented classes and enhancing overall model robustness.
@@ Acknowledgement
This research was supported by the Fundamental Research Funds for the Central Universities (No.2682025CX082), the National Natural Science Foundation of China (Nos. 62176221, 61572407), and Sichuan Province Science and Technology Support Program (Nos. 2024NSFTD0036, 2024ZHCG0166).